Q. Derive 3rd law of newton using 2nd law. What value does the slope of static friction vs applied force have apparently?

Grade:11

1 Answers

Arun
5 years ago
let the net momentum of a system of two bodies of mass m1 and m2 be p where p=p1+p2.from the second law net external force on the body =dp1/dt+dp2/dt. if the total momentum in any direction is constant then dp1/dt+dp2/dt=0(derivative of a constant is zero.)let the velocities of the bodies change from u1 to v1 and u2 to v2 in time t due to their mutual interaction. then m1(v1-u1)/t+m2(v2-u2)/t=0 (since dp/dt is time rate of change of momentum.) again we know (v-u)/t=acceleration. so m1f1+m2f2=0. again from the second law we get force =mass*acceleration. so f1+f2=0 and so f1= -f2. which shows thr force applied by the first body on the second is equal in magnitude but opposite in direction to the force applied by the second on the first. this is stated in the third law.thus it is derieved from the second law.